OpenStudy (anonymous):

solve for y in the first equation. x + 3y = 6 2x – y = 10

OpenStudy (anonymous):

x - 3y + 2x - 7y = - 6 + 10 10y = 3x - 4 y = 0.3x - 0.4 x - 3(0.3x - 0.4) = - 6 x - 0.9x + 1.2 = - 6 0.1x = - 7.2 x = - 72 y = 0.3(- 72) - 0.4 y = - 21.6 - .4 y = - 22 Answer: x = - 72, y = - 22

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Sorry, stepped away

OpenStudy (anonymous):

My answer choices are ..

OpenStudy (anonymous):

1/3x+2 -1/3x+6 -x+2 -1/3+2